Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Q

Expert replies
by punitkaur » Mon Nov 02, 2009 8:50 am
. What is the probability for a family with three children to have a boy and two girls (assuming the probability of having a boy or a girl is equal)?

a) 1/8
b) 1/4
c) 1/2
d) 3/8
e) 5/8

My answer is 1/8 but the answer given is 3/8
Join the discussion
Source: — Problem Solving |

by jzdchou » Mon Nov 02, 2009 10:20 am
It is like flipping a coin 3 times and getting 8 different total outcomes.

1/2 * 1/2 * 1/2 = 1/8

Then I just list the possibilities
BBB
BGG
BBG
GBB
GGB
GBG
BGG
GGG

2G and 1B = GGB, GBG, BGG = 3/8
Join the discussion

by punitkaur » Mon Nov 02, 2009 10:26 am
Thanks jzdchou. I did a silly mistake while solving the problem, i took all 3 cases GGB, GBG, BGG to be the same and hence got 1.

Did not realize that sequence does matter.
Join the discussion

by jzdchou » Mon Nov 02, 2009 10:30 am
Or once you get total of 8 combinations....

instead of doing the listing, think of how many ways can 2G and 1B be arranged/ordered.

Isn't it simply 3!/2! = 3*2*1/2*1 = 3?
Join the discussion

ans

by crackthetest » Sat Nov 07, 2009 8:09 pm
How to know if you need to use the permutation here, so that order really matters. I see the question just asks for one boy & 2 girls but not how they are arranged.

Pl. clarify.
Join the discussion

Re: ans

by Stuart@KaplanGMAT » Mon Nov 09, 2009 10:29 am
crackthetest wrote:How to know if you need to use the permutation here, so that order really matters. I see the question just asks for one boy & 2 girls but not how they are arranged.

Pl. clarify.
Since the question doesn't mention a specific order, all orders are acceptable. Since all orders are acceptable, we have to take them all into account.

Identifying this as a pseudo-coin flip question makes life much easier, especially if you know the coin flip formula:

Prob(k results out of n flips) = (nCk)/2^n

In this question, we want 2 boys out of 3 kids, so let's let n=3 and k=2:

(3!/2!1!)/(2^3) = 3/8

Note that we could have looked at the question from the other side: we want 1 girl out of 3 kids, so let's let n=3 and k=1:

(3!/1!2!)/(2^3) = 3/8

Any question with a 50/50 probability can be solved using the above formula.

For more info about coin flip and pseudo-coin flip questions, check out this thread:

https://www.beatthegmat.com/coin-flip-qu ... 17911.html
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

Thanks

by crackthetest » Sat Nov 14, 2009 9:15 pm
Thanks very much Stuart, your link to pseudo-coin is very helpful.
Join the discussion